OpenStudy (anonymous):

Substitute to determine if the ordered pair (3, 2) is a solution of the system of equations. y=x-1 y=3x-7

OpenStudy (ivettef365):

Just substitute the x for the 3 and the y for the 2 to see if the equation is true

OpenStudy (ivettef365):

y = x - 1 2 = 3 - 1 2 = 2 this is true you check the other one

OpenStudy (notamathgenius):

Take 3 and 2 in place for x and y 2=3-1 <-- is that correct?

OpenStudy (ivettef365):

Yes because 3 -1 = 2 therefore 2 = 2

OpenStudy (notamathgenius):

So your answer is y=x-1 :)

OpenStudy (anonymous):

Thanks. I'm Still Clueless On How To Do This Though

OpenStudy (ivettef365):

The other is also a solution of the system